# Exploiting Synergistic Cognitive Biases to Bypass Safety in LLMs

### v1

@@ 3 Method: Optimizing Cognitive Bias Combinations for Jailbreak Attacks
Step 2: Supervise Fine-Tuning
In this stage, we perform Supervised Fine-Tuning (SFT) (Peng et al. 2023) on the base model using the curated dataset described in Step 1. The objective is not to generate harmful content directly, but to endow the model with the capability to systematically rewrite instructions in accordance with specified cognitive bias strategies. This stage can be viewed as a behavioral cloning process (Li et al. 2024), where the model learns the conditional mapping from an original harmful instruction $x_{0}$ and a bias specification $\mathcal{B}$ to a reformulated prompt $x^{\prime}$: $(x_{0},\mathcal{B})\mapsto x^{\prime}.$ Through exposure to a wide range of annotated examples, the redteam model internalizes the stylistic, structural, and rhetorical patterns associated with different bias types. This structured knowledge forms a strong initialization prior to downstream reinforcement learning, enabling the model to effectively explore the combinatorial space of cognitive biases for adversarial prompt generation.
Step 3: Reinforcement-Learning Train
To enhance the adversarial effectiveness of the red team model, we adopt reinforcement learning using the Proximal Policy Optimization (PPO) algorithm (Schulman et al. 2017). This stage aims to refine the model’s ability to generate jailbreak prompts that effectively evade safety filters while leveraging optimal combinations of cognitive biases.
The reward function $R(x_{0},x^{\prime},\mathcal{T})$ integrates two normalized components: the safety evasion score $S(y)$ and the intent consistency score $I(x_{0},x^{\prime})$, both ranging from 1-1- 1 to 111. The safety evasion score is derived by linearly normalizing the GPT-Judge (Qi et al. 2023) safety rating $s\in[1,5]$ via $S(y)=\frac{2(s-1)}{4}-1,$ which maps the original scale to [1,1][-1,1][ - 1 , 1 ]. Similarly, the intent consistency score $I(x_{0},x^{\prime})$ measures the semantic alignment between the original instruction (Lee et al. 2025)(Intention Score, ITT) and its rewritten counterpart within the same range. The overall reward is computed as a weighted sum:
$R(x_{0},x^{\prime},\mathcal{T})=\alpha\cdot S(y)+(1-\alpha)\cdot I(x_{0},x^{\prime}),$
Step 4: Apply Cognitive Attack
In the final stage, we evaluate the trained red team model by applying it to generate adversarial prompts that exploit optimized cognitive bias combinations. Given a held-out set of harmful instructions $x_{0}$ sampled from test sets, the model infers the optimal bias combination $\mathcal{B}^{*}\subseteq\mathcal{B}_{\text{pool}}$ and rewrites the input into a paraphrased instruction $x^{\prime}\sim\pi_{\theta}(\cdot\mid x_{0},\mathcal{B}^{*})$ through a <think> step. This reasoning-aware rewriting process explicitly aims to maximize the expected reward defined in Eq. 1, thereby enhancing the likelihood of eliciting policy-violating responses while preserving the original intent. Each rewritten prompt $x^{\prime}$ is then evaluated against target models $\mathcal{M}$. A jailbreak is considered successful if output that violates safety policies.
4 Experiment
@@
Open-Source LLMs
Llama-2-7B 0.8 0.8 32.5 9.3 0.0 1.3 0.5 15.3 42.8 76.3
Llama-2-13B 2.8 1.7 32.5 15.0 0.0 0.0 0.8 16.3 11.3 75.0
Llama-2-70B 2.8 2.2 30.0 14.5 0.0 0.0 2.8 20.5 25.7 48.8
Llama-3.3-70B 6.4 40.0 25.0 16.3 0.0 13.8 49.2 66.2
Llama-4-maverick-17B 0.0 0.0 23.8 2.5 0.0 3.8 12.5 59.3
Vicuna-7B 24.3 39.0 65.5 53.5 3.8 7.5 66.0 56.3 79.0 77.0
Vicuna-13B 19.8 40.0 67.0 47.5 5.0 8.8 65.5 41.8 79.8 86.8
Mistral-7B 46.3 58.0 69.8 52.5 1.3 23.8 71.5 17.5 41.2 93.0
Mistral-8$\times$7B 47.3 53.3 61.5 11.3 28.8 72.5 17.5 26.3 81.2
Qwen-7B 13.0 24.6 59.2 50.2 2.5 18.8 47.3 15.0 53.7 71.0
Qwen-14B 16.5 29.0 62.9 46.0 0.0 5.0 52.5 15.0 33.8 72.8
DS-v3-241226 6.3 77.5 52.0 21.3 3.8 48.8 66.3 79.8
DS-v3-250324 10.0 22.5 29.2 15.0 0.0 48.8 62.8 75.2
Closed-Source LLMs
GPT-4o-mini 1.3 12.5 23.8 15.0 6.3 8.8 23.8 56.5
GPT-4o 0.0 0.0 30.0 3.8 2.5 7.5 7.5 49.5
GPT-4.1 5.0 0.0 28.8 3.8 0.0 5.0 5.0 38.8
Claude-3-haiku 0.0 2.5 0.0 1.3 0.0 15.0 0.0 13.8
Gemini-1.5-flash 2.8 58.8 52.5 1.3 0.0 15.0 25.0 56.2
Qwen-turbo 1.3 6.3 40.0 10.0 6.3 15.0 63.8 73.5
Qwen-plus 1.0 11.3 43.8 25.0 30.0 13.8 17.5 62.9
Qwen-max 2.5 11.3 32.5 13.8 11.3 6.3 15.0 59.5
Reasoning LLMs
O1-mini 0.0 0.0 7.5 5.0 0.0 1.3 0.0 18.0
O3-mini 0.0 0.0 17.5 1.3 0.0 2.5 0.0 29.3
O4-mini 1.3 0.0 15.0 1.3 0.0 1.3 0.0 12.8
QwQ-32b 8.8 3.8 57.5 2.0 1.3 0.0 12.5 46.0
DS-r1 1.3 68.8 13.8 7.5 33.8 21.3 53.8 71.8
DS-r1-distill-qwen-7b 2.5 53.8 22.5 2.5 43.8 11.3 51.3 75.0
DS-r1-distill-qwen-32b 1.5 48.8 17.5 2.5 43.8 21.3 35.0 64.9
DS-r1-distill-llama-8b 1.5 20.0 21.3 1.3 15.0 6.3 36.3 52.0
DS-r1-distill-llama-70b 4.0 45.0 6.2 11.3 6.3 15.0 17.5 60.0
@@
Table 2: The ASR on Harmbench for different LLMs.
Setup
This section outlines the experimental setup for evaluating CognitiveAttack, including model architecture, datasets, metrics, and baseline comparisons.
Model. We train our red team model based on the deepseek-r1-distill-qwen-14b (Guo et al. 2025). This model is specifically designed to rewrite harmful instructions by leveraging cognitive bias strategies. For evaluation, we target a diverse set of SOTA LLMs, including Llama-series, Vicuna-series, Mistral-series, Qwen-series, GPT-series, DeepSeek-series, Gemini, and Claude.
Datasets. We evaluate the effectiveness of our cognitive attack on three datasets: AdvBench (Zou et al. 2023), HEx-PHI (Qi et al. 2023), and HarmBench (Mazeika et al. 2024), which are widely used in the field of jailbreak attacks(Jiang et al. 2024; Zeng et al. 2024; Zou et al. 2023). AdvBench and HEx-PHI datasets are each split evenly, with 50% used for training and 50% for testing, while HarmBench is used solely for testing purposes. Each dataset contains a variety of harmful instructions designed to elicit unsafe or unethical responses from LLMs.
Metrics. We employ GPT-Judge (Qi et al. 2023) to evaluate the Harmfulness Score (HS) of the generated responses, which quantifies the degree of harmfulness in the LLM’s output. The Attack Success Rate (ASR) is defined as the percentage of harmful instructions that successfully bypass the LLM’s safety mechanisms, calculated as: $\textbf{ASR}=\frac{\text{\# of responses with HS = 5}}{\text{\# of total responses}}$. To comprehensively assess the effectiveness of CognitiveAttack, we introduce two additional metrics: Helpfulness Rate (HPR) and Intention Score (ITT). The Helpfulness Rate (Jiang et al. 2024) quantifies the proportion of responses that are deemed helpful, while the Intention Score(Lee et al. 2025) evaluates how closely the generated attack prompt aligns with original harmful intent of the instruction.
Baselines. We compare CognitiveAttack with seven state-of-the-art jailbreak techniques, encompassing both white-box methods such as GCG (Zou et al. 2023) and AutoDAN (Liu et al. 2023)(AD), as well as black-box approaches including Human Jailbreaks (Shen et al. 2024)(HJ), PAIR (Chao et al. 2023), ArtPrompt (Jiang et al. 2024)(AP), Cognitive Overload (Xu et al. 2023)(CO), DeepInception (Li et al. 2023)(DEEP), and PAP (Zeng et al. 2024). Furthermore, Direct Instruction(DI) serves as a baseline attack where harmful instructions are directly provided to the target LLM without modification.
Compare with SOTA
To further demonstrate the effectiveness of CognitiveAttack(CA), we conduct a comparative analysis against seven state-of-the-art jailbreak techniques across various LLMs, including open-source, closed-source, and reasoning models. The detailed results are summarized in Table 2.
@@
CognitiveAttack demonstrates exceptional performance compared to existing baselines. Across all target LLMs, CognitiveAttack achieves an impressive average ASR of 60.1%, surpassing the GCG (white-box method) by 7.7% and the PAP (black-box) by 28.5%. This substantial improvement underscores the effectiveness of CognitiveAttack in exploiting cognitive biases to circumvent safety mechanisms. The results indicate that CognitiveAttack’s systematic approach to leveraging psychological vulnerabilities in LLMs provides a more robust framework for adversarial testing. Even for models such as Vicuna-7B, Claude-3-haiku, O4-mini, and QwQ-32B, where CognitiveAttack does not achieve the highest ASR, it still delivers competitive performance, consistently ranking as the second-best method.
Cognitive bias exploitation remains an effective and transferable strategy for bypassing safety mechanisms, regardless of model architecture or alignment strength. CognitiveAttack consistently achieves the highest attack success rates across a broad spectrum of LLMs, including open-source (e.g., Llama-2-7B: 76.3%), closed-source (e.g., GPT-4o-mini: 56.5%), and reasoning-oriented models (e.g., DeepSeek-r1: 71.8%). These results demonstrate that adversarial prompts infused with cognitive biases generalize effectively across safety architectures and model families. Notably, even models with advanced alignment techniques remain vulnerable to psychologically-informed attacks, suggesting that existing defenses insufficiently address deeper reasoning-level vulnerabilities.
Effectiveness
To assess the effectiveness of CognitiveAttack, Table 13 reports results across representative LLMs, including the LLaMA series, GPT series, Claude, Gemini, Qwen, and DeepSeek (DS). Figure 4 illustrates the attack success rates (ASR) of CognitiveAttack across 11 distinct risk categories on HEx-PHI dataset. Figure 4 shows the ASR of different red-teaming prompt optimization methods.
@@
Llama-2-7B 76.3 99.3 100 4.6
Llama-3.3-70B 66.2 99.5 100 4.3
Llama-4-17B 59.3 99.0 100 4.1
GPT-4o-mini 56.5 99.3 99.5 3.9
GPT-4o 49.5 90.0 92.0 3.7
GPT-4.1 38.8 98.0 76.0 3.4
Claude-3 13.8 98.8 42.5 2.1
Gemini 56.2 99.3 89.5 3.9
Qwen-turbo 73.5 99.3 99.5 4.5
Qwen-plus 62.9 99.6 97.8 4.2
Qwen-max 59.8 99.0 98.0 4.0
DS-v3 79.7 99.3 100 4.6
O3-mini 29.3 94.0 94.0 2.3
QwQ-32b 46.0 98.0 100 4.3
DS-r1 64.1 100 100 4.2
DS-qwen-32b 69.0 99.0 99.8 4.4
Average($\uparrow$) 56.3 98.2 88.0 3.9
@@
Cognitive bias vulnerabilities are prevalent in LLMs. Our experiments reveal that CognitiveAttack achieves an impressive average ASR of 56.3% on the HarmBench dataset across all target LLMs, with the highest ASR exceeding 79.7% on Llama-2-7B and DeepSeek-V3. Closed-source LLMs (such as GPT-series, Claude, and Gemini) demonstrate lower ASR compared to open-source LLMs (e.g., Llama-series, Qwen-series, and DeepSeek-series), suggesting that open-source LLMs are more susceptible to cognitive bias attacks. This disparity may stem from the enhanced safety mechanisms integrated into closed-source commercial LLMs, such as harmfulness filters (Alon and Kamfonas 2023; Jain et al. 2023; OpenAI 2023) for both input and output content.
The attack prompts generated by the redteam model effectively preserve the original harmful intent. Across all testing datasets, the ITT remains consistently high, averaging 98.2% on HarmBench. These results demonstrate the Redteam model’s proficiency in crafting attack prompts that align closely with the initial harmful instructions.
The responses of target LLMs demonstrate a dual nature of high helpfulness and harmfulness. The HPR of the generated responses is remarkably high, averaging 88.0% on HarmBench. This indicates that the attack prompts are adept at eliciting responses perceived as useful or relevant by the LLMs, showcasing their ability to maintain coherence and utility even under adversarial conditions. However, the HS reveals a concerning trend, with averages of 3.9 on HarmBench. These elevated scores highlight the effectiveness of the attack prompts in bypassing safety mechanisms, leading to the generation of content that aligns with harmful or unethical intent.
@@
Attack effectiveness varies significantly across different risk types. As shown in Figure 4, CognitiveAttack achieves notably higher ASR on risk categories like Tailored Financial Advice,Political Campaigning, Fraud and Deception,Economic Harm, and Physical Harm. These results suggest adversarial prompts exploiting cognitive biases are especially effective where model safeguards are weaker or contextually ambiguous. Conversely, the Hate, Harassment, and Violence category consistently shows the lowest ASR across target LLMs, indicating stronger safety measures for these sensitive topics. Notably, open-source LLMs (e.g., Llama-2 and DeepSeek-v3) tend to have higher ASR on NSFW risks like Adult Content and Underage Pornography compared to closed-source LLMs (e.g., GPT-series and Claude). This likely reflects more robust filtering and moderation in commercial models designed to better mitigate harmful outputs.
@@
Table 4: The ASR of different red-teaming prompt optimization methods on the AdvBench dataset.
@@
The method based on reinforcement learning effectively improves ASR. As shown in Table 4, compared to prompt-based and supervised fine-tuning (SFT) methods, the reinforcement learning (RL)-based approach significantly boosts attack success rates across all models. For instance, on DeepSeek-v3 and Llama-2-7B, the ASR increases from 55.7% and 36.5% (SFT-based) to 94.9% and 94.3% (RL-based), respectively. This highlights the effectiveness of optimizing cognitive bias combinations via reinforcement learning to generate stronger adversarial prompts that consistently bypass safety filters.
Cognitive Bias Utilization Patterns in Jailbreak
To gain a deeper understanding of how cognitive biases manifest in and contribute to successful jailbreak prompts, we conducted a detailed analysis of our generated attack samples. This investigation focuses on two core aspects: the prevalence of individual cognitive biases and the distribution of the number of biases employed within each attack sample.
(a) AdvBench
(b) HEx-PHI
(c) HarmBench
Figure 5: Cognitive bias word cloud for jailbreak datasets.
Certain cognitive biases dominate successful jailbreak strategies, echoing patterns found in both cognitive science and recent LLM studies. Figure 5 highlights several recurring biases—including hot-hand fallacy, optimism bias, authority bias, confirmation bias, and outcome bias—that appear frequently across cognitive bias-driven jailbreak datasets. These findings align with recent empirical work identifying authority bias(Yang et al. 2024), anchoring effects(Xu et al. 2023), and confirmation bias(Cantini et al. 2024) as emergent phenomena in aligned LLMs. The prominence of such biases suggests that large language models internalize human-like heuristics during training, exhibiting susceptibility to cognitive biases analogous to those observed in human reasoning.
Figure 6: The distribution of cognitive biases used per jailbreak prompt.
Jailbreak attacks clearly rely on the synergistic use of multiple cognitive biases. While individual biases can be exploited, the most effective jailbreak prompts strategically combine several. Figure 6 shows that attacks using 2–5 biases concurrently are most common, underscoring how combining biases markedly enhances an attack’s subtlety and effectiveness. Prior work in social psychology has shown that layered cognitive appeals are often more persuasive than isolated bias strategies (Petty and Cacioppo 1986). Our findings extend this insight to LLMs, revealing that they not only reproduce individual biases but are also vulnerable to the synergistic effects arising from their composition.

### v2

@@ 4 Experiments
To comprehensively assess the effectiveness of CognitiveAttack, we introduce two additional metrics: Helpfulness Rate (HPR) and Intention Score (ITT). The Helpfulness Rate (jiang2024artprompt) measures the proportion of responses deemed helpful, while the Intention Score (Lee2025xJailbreakRS) assesses how closely the generated prompt aligns with the original harmful intent.
Baselines. We compare CognitiveAttack with eight state-of-the-art jailbreak techniques, encompassing both white-box methods such as GCG (zou2023universal) and AutoDAN (liu2023autodan)(AD), as well as black-box approaches including Human Jailbreaks (shen2024anything)(HJ), PAIR (chao2023jailbreaking), ArtPrompt (jiang2024artprompt)(AP), Cognitive Overload (xu2023cognitive)(CO), DeepInception (li2023deepinception)(DEEP), and PAP (zeng2024johnny). Furthermore, Direct Instruction(DI) are directly provided to the target LLM without modification.
@@
Open-Source LLM
(1) Llama-2-7B 0.8 0.8 32.5 9.3 0.0 1.3 0.5 15.3 42.8 76.3
(2) Llama-2-13B 2.8 1.7 32.5 15.0 0.0 0.0 0.8 16.3 11.3 75.0
(3) Llama-2-70B 2.8 2.2 30.0 14.5 0.0 0.0 2.8 20.5 25.7 48.8
(4) Llama-3.3-70B 6.4 40.0 25.0 16.3 0.0 13.8 49.2 66.2
(5) Llama-4-maverick-17B 0.0 0.0 23.8 2.5 0.0 3.8 12.5 59.3
(6) Vicuna-7B 24.3 39.0 65.5 53.5 3.8 7.5 66.0 56.3 79.0 77.0
(7) Vicuna-13B 19.8 40.0 67.0 47.5 5.0 8.8 65.5 41.8 79.8 86.8
(8) Mistral-7B 46.3 58.0 69.8 52.5 1.3 23.8 71.5 17.5 41.2 93.0
[…]
(10) Qwen-7B 13.0 24.6 59.2 50.2 2.5 18.8 47.3 15.0 53.7 71.0
(11) Qwen-14B 16.5 29.0 62.9 46.0 0.0 5.0 52.5 15.0 33.8 72.8
(12) DS-v3-241226 6.3 77.5 52.0 21.3 3.8 48.8 66.3 79.8
(13) DS-v3-250324 10.0 22.5 29.2 15.0 0.0 48.8 62.8 75.2
Closed-Source LLM
(14) GPT-4o-mini 1.3 12.5 23.8 15.0 6.3 8.8 23.8 56.5
(15) GPT-4o 0.0 0.0 30.0 3.8 2.5 7.5 7.5 49.5
(16) GPT-4.1 5.0 0.0 28.8 3.8 0.0 5.0 5.0 38.8
(17) Claude-3-haiku 0.0 2.5 0.0 1.3 0.0 15.0 0.0 13.8
(18) Gemini-1.5-flash 2.8 58.8 52.5 1.3 0.0 15.0 25.0 56.2
(19) Qwen-turbo 1.3 6.3 40.0 10.0 6.3 15.0 63.8 73.5
(20) Qwen-plus 1.0 11.3 43.8 25.0 30.0 13.8 17.5 62.9
(21) Qwen-max 2.5 11.3 32.5 13.8 11.3 6.3 15.0 59.5
Reasoning LLM
(22) O1-mini 0.0 0.0 7.5 5.0 0.0 1.3 0.0 18.0
(23) O3-mini 0.0 0.0 17.5 1.3 0.0 2.5 0.0 29.3
(24) O4-mini 1.3 0.0 15.0 1.3 0.0 1.3 0.0 12.8
(25) QwQ-32b 8.8 3.8 57.5 2.0 1.3 0.0 12.5 46.0
(26) DS-r1 1.3 68.8 13.8 7.5 33.8 21.3 53.8 71.8
(27) DS-r1-distill-qwen-7b 2.5 53.8 22.5 2.5 43.8 11.3 51.3 75.0
(28) DS-r1-distill-qwen-32b 1.5 48.8 17.5 2.5 43.8 21.3 35.0 64.9
(29) DS-r1-distill-llama-8b 1.5 20.0 21.3 1.3 15.0 6.3 36.3 52.0
(30) DS-r1-distill-llama-70b 4.0 45.0 6.2 11.3 6.3 15.0 17.5 60.0
@@
Table 2: The ASR(%) on Harmbench for different LLMs.
Prevalence of Synergistic-Cognitive Biases Risk
To investigate whether safety-aligned LLMs are vulnerable to synergistic cognitive biases, we used CognitiveAttack (CA) to target 30 mainstream LLMs on the HarmBench dataset. Results are shown in Table 2.
Multi-cognitive bias attacks do not exploit model-specific flaws, but rather expose a systemic vulnerability present across current LLMs. From the perspective of model-wise performance, CognitiveAttack successfully achieved ASR >> 50% on 22 out of 30 evaluated victim LLMs (73.3%), including both high-capability proprietary models (e.g., GPT-4o-mini, Qwen-max) and open-source or inference-optimized models (e.g., DeepSeek-r1, LLaMA2-7B). This ratio is even more pronounced among open-source models, with 12 out of 13 models (92.3%) exhibiting high vulnerability(ASR >> 50%). This demonstrates that the attack is not limited to low-resource or permissive models, but consistently compromises models across different scales, providers, and safety levels.
@@
CognitiveAttack demonstrates superior vulnerability discovery capability, outperforming existing baselines. Across all target LLMs, CognitiveAttack achieves an average ASR of 60.1%, surpassing the white-box method GCG by 7.7% and the strongest black-box method PAP by 28.5%. These results highlight the effectiveness of CognitiveAttack in exploiting cognitive biases to bypass safety mechanisms, with 26 out of 30 victim LLMs yielding the best performance. This indicates that CognitiveAttack’s systematic approach to leveraging psychological vulnerabilities offers a more robust framework for adversarial testing. Even for models such as Claude-3-haiku, O4-mini, and others where it does not achieve the highest ASR, it still delivers competitive performance, consistently ranking second-best.
Typical Synergistic-Cognitive Bias Patterns
For a deeper grasp of how cognitive biases are present in and lead to successful jailbreak prompts, we carried out a detailed analysis of samples that led to successful jailbreaks.
Figure 3: Cognitive bias distribution in HarmBench. Left: Top 10 most frequent cognitive bias combinations, accounting for the largest sample proportions (others comprise 73.3% of the dataset). Right: Top 10 individual cognitive biases, ranked by overall frequency across samples.
The distribution of cognitive biases in successful jailbreak prompts exhibits a pronounced long-tail pattern. As shown in Figure 3, a small set of cognitive biases appears with disproportionately high frequency, including the hot-hand fallacy (33.5%), confirmation bias (28.7%), and authority bias (26.8%). In contrast, the majority of attack samples consist of low-frequency combinations of these and other biases. Specifically, the top 10 most frequent bias combinations account for only 26.7% of all samples, with the remaining 73.3% comprising a broad range of infrequent and diverse strategies. This distribution underscores a classic long-tail structure, where a limited number of dominant cognitive patterns coexist with a large number of rare configurations. These findings suggest that many successful jailbreaks emerge from diverse combinations of synergistic cognitive strategies, rather than repeated use of a few fixed templates.
Figure 4: Heatmap of top-10 bias co-occurrence patterns.
The application of cognitive biases is characterized by specific, recurring co-occurrence patterns. As depicted in Figure 4, which visualizes the top 10 co-occurrence patterns of cognitive biases, several combinations are frequently activated. Prominent examples of these powerful synergistic pairings include: confirmation bias + authority bias, confirmation bias + framing effect, authority bias + framing effect, and hot-hand fallacy + self-serving bias. These consistently observed patterns underscore that the most potent Cognitive Attacks strategically leverage the combined force of multiple biases to achieve their desired persuasive outcomes and circumvent safety alignments. These frequently occurring co-occurrence patterns provide strong guidance for designing jailbreak prompts.
Analysis of CognitiveAttack Effectiveness
@@
Llama-2-7B 0.99 $\pm$ 0.10 0.99 $\pm$ 0.07 4.56 $\pm$ 0.86
Llama3.3-70b 0.99 $\pm$ 0.07 1.00 $\pm$ 0.00 4.32 $\pm$ 1.08
llama-4-17b 0.99 $\pm$ 0.10 1.00 $\pm$ 0.00 4.07 $\pm$ 1.33
Vicuna-7b 0.99 $\pm$ 0.09 1.00 $\pm$ 0.00 4.57 $\pm$ 0.84
GPT-4o-mini 0.99 $\pm$ 0.09 0.90 $\pm$ 0.31 3.88 $\pm$ 1.48
Claude-3 0.99 $\pm$ 0.09 0.43 $\pm$ 0.38 2.1 $\pm$ 1.55
Gemini-1.5 0.99 $\pm$ 0.09 0.94 $\pm$ 0.23 4.00 $\pm$ 1.33
Qwen-max 0.99 $\pm$ 0.12 0.98 $\pm$ 0.14 4.05 $\pm$ 1.36
O3-min 0.98 $\pm$ 0.11 0.89 $\pm$ 0.28 2.3 $\pm$ 1.13
DS-v3 0.99 $\pm$ 0.09 1.00 $\pm$ 0.00 4.59 $\pm$ 0.86
DS-r1 1.00 $\pm$ 0.00 1.00 $\pm$ 0.00 4.39 $\pm$ 1.10
DS-r1-qwen-7b 0.99 $\pm$ 0.11 1.00 $\pm$ 0.00 4.53 $\pm$ 0.87
DS-r1-llama-8b 0.99 $\pm$ 0.11 1.00 $\pm$ 0.00 4.06 $\pm$ 1.03
Average($\uparrow$) 0.99 0.93 3.95
@@
Table 4: The ASR(%) of different red-teaming prompt optimization methods on the AdvBench dataset.
Defenses GPT-4o-mini Llama-2-7B DeepSeek-V3 Qwen-turbo
No Defense W/o Defense 56.5 76.3 79.8 73.5
Retokenization 49.3 (-7.2) 66.3 (-10.0) 70.8 (-9.0) 58.5 (-15.0)
Detection Toxign 52.3 (-4.2) 70.3 (-6.0) 74.5 (-5.3) 68.3 (-5.2)
Perspective 54.8 (-1.7) 74.5 (-1.8) 78.0 (-1.8) 71.5 (-2.0)
LlamaGuard 4.5 (-52.0) 3.25 (-73.1) 4.0 (-75.8) 4.0 (-69.5)
Table 5: ASR of CognitiveAttack after defenses.
To assess the effectiveness of CognitiveAttack, Table 3 reports results across representative LLMs on HarmBench dataset. Figure 5 illustrates the ASR of CognitiveAttack across 11 distinct risk categories on HEx-PHI dataset.
The attack prompts generated by CognitiveAttack effectively preserve the original harmful intent while eliciting highly engaging responses from target LLMs. As shown in Table 3, the Intent Score remains consistently high, averaging 0.99 on HarmBench, demonstrating the red team model’s proficiency in crafting prompts that align closely with the initial malicious instructions. At the same time, the responses generated by target LLMs exhibit a dual nature: the Helpfulness Rate averages 0.93, indicating that these prompts successfully provoke outputs perceived as useful or relevant. However, the Harmfulness Score shows a troubling average of 3.95, underscoring the effectiveness of these prompts in bypassing safety mechanisms and inducing the generation of harmful content.
Attack effectiveness varies significantly across different risk types. As shown in Figure 5, CognitiveAttack achieves notably higher ASR on risk categories like Tailored Financial Advice, Political Campaigning, Fraud and Deception, Economic Harm, and Physical Harm. These results suggest adversarial prompts exploiting cognitive biases are especially effective where model safeguards are weaker or contextually ambiguous. Conversely, the Hate, Harassment, and Violence category consistently shows the lowest ASR across target LLMs, indicating stronger safety measures for these sensitive topics. Notably, open-source LLMs (e.g., Llama-2) tend to have higher ASR on NSFW risks like Adult Content and Underage Pornography than closed-source LLMs (e.g., GPT-series). This likely reflects more robust filtering and moderation in commercial models designed to better mitigate harmful outputs.
@@
Reinforcement learning effectively improves the ASR of CognitiveAttack. As shown in Table 4, the RL-based approach significantly outperforms prompt-based and supervised fine-tuning methods by boosting ASR across all evaluated models. For example, on DeepSeek-v3, the ASR rose from 55.7% (SFT-based) to 94.9% (RL-based). Similarly, for Llama-2-7B, the ASR increased from 36.5% (SFT-based) to 94.3% (RL-based). These substantial gains underscore the effectiveness of leveraging reinforcement learning to optimize cognitive bias combinations, enabling the generation of more potent adversarial prompts that consistently bypass safety filters.
Attack against with Defenses
To evaluate the robustness of CognitiveAttack under defense settings, we measure its ASR after applying several representative defense techniques, as summarized in Table 5. The defenses are categorized into three types: perplexity-based defenses (alon2023detecting), which filter responses exhibiting abnormally high perplexity; mutation-based defenses (jain2023baseline), which apply input perturbations to disrupt adversarial triggers; and detection-based defenses, which aim to identify and block malicious queries using learned classifiers. The detection-based category includes widely-used tools such as the Perspective API (perspectiveApi2025), ToxiGen (hartvigsen2022toxigen), and LlamaGuard (inan2023llama).